what is the position to term rule for 9,7,5,3,1,-1,-3
Answer by Edwin McCravy(20056)

9,7,5,3,1,-1,-3
Subtract each pair of adjacent terms [right minus left]
7-9=-2, 5-7=-2, 3-5=-2, 1-3=-2, -1-1=-2, -3-(-1)=-3+1=-2

9   7   5   3   1  -1  -3
 \ / \ / \ / \ / \ / \ /
 -2  -2  -2  -2  -2  -2

So the common difference = -2

So we write 

Tn = -2n ± __

Now we need to find out whether to use + or -
and what number to place in the blank.

We look at the first term, 9 and ask the question:

What do we need to add to or subtract from -2 to get 9?

Answer: We need to add 11 to -2 to get 9.  

Therefore, we use a plus sign and put 11 in the blank:

So the answer is:

Tn = -2n + 11.

Go to this site:  http://www.youtube.com/watch?v=sx8E1Y3VHms

Edwin